OpenStudy (anonymous):

Identify the oblique asymptote of f(x) = quantity x squared minus 4 x plus 8 over quantity x plus 2. y = 0 y = x - 2 <------ MY ANSWER y = x - 6 No oblique asymptote Can you please check this for me? @jim_thompson5910

OpenStudy (anonymous):

oh here we go with the "overs" again

OpenStudy (anonymous):

LOL you're funny! xD

OpenStudy (anonymous):

\[\frac{x^2-4x+6}{x+2}\]

OpenStudy (anonymous):

is that the right thing?

OpenStudy (anonymous):

if so unfortunately your answer is wrong

OpenStudy (anonymous):

+8 instead of +6 for the numerator

OpenStudy (anonymous):

still wrong

OpenStudy (anonymous):

oh :/ ughhhh

OpenStudy (anonymous):

\[\frac{x^2-4x+8}{x+2}\] you know how to do synthetic division? it is real easy here

OpenStudy (anonymous):

no

OpenStudy (anonymous):

damn you know polynomial long division?

OpenStudy (anonymous):

no lol i'm STOOOOOPID! if you haven't noticed by now

OpenStudy (anonymous):

lol stoopid is as stoopid does you know wolfram alpha?

OpenStudy (anonymous):

http://www.wolframalpha.com/input/?i=%28x^2-4x%2B6%29%2F%28x%2B2%29 you gotta divide when you divide you get a quotient and a remainder ignore the remainder, the quotation is the slant asymptote

OpenStudy (anonymous):

there is no way to do this without doing the division \[\frac{x^2-4x+8}{x+2}=x-6+\frac{18}{x+2}\]

OpenStudy (anonymous):

the slant asymptote is \(y=x-6\)

OpenStudy (anonymous):

thank you senor!! :P

OpenStudy (anonymous):

de nada madam